OpenStudy (anonymous):

Can I please get some help? a+8B=16 a+2b=-2

OpenStudy (pfenn1):

Subtract the 2nd equation from the first equation and solve for b.\[+a+8b=16\]\[-a-2b=2\]\[=>(8-2)b=16+2\]\[=> 6b=18\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

okay, so would it be one solution? Infinitely many solutions? or no solution?

OpenStudy (anonymous):

You would have a solution for a and one for b. So 2 solutions :)

OpenStudy (anonymous):

I came up with (-8,3) am I right?

OpenStudy (jiteshmeghwal9):

First find the value of 'a' from any of the equation. Then put the value of 'a' in any of the equation.Similarly,find the value of 'b' and put it into one of the equation and solve it.and,then u will get the value of 'a' and 'b'.